1.    A radioactive sample with a half-life of  3hrs contains 25-mCi radioactivity. The radioactivity of the sample after 9 days days would be?

2.1.    A sample of material contains 1 microgram of iodine-131. Note that, iodine-131 plays a major role as a radioactive isotope present in nuclear and it a major contributor to the health hazards when released into the atmosphere during an accident. Iodine-131 has a half-life of 8.02 days.

 

a)The number of iodine-131 atoms initially present.

b)The activity of the iodine-131 in curies.

c)The number of iodine-131 atoms that will remain in 50 days.

d)The time it will take for the activity to reach 0.1 mCi.

 


1
Expert's answer
2021-10-10T16:02:12-0400

1. 9 days = 9"\times"24 = 216 hours, which "\frac{216}{3} =" 72 half-life periods

The activity will be "\frac{25}{2^{72}} = 5.29 \times10^{-21} mCi"

2.

a) 1 microgram is "10^{-6}" g; the number of atoms is "\frac{10^{-6}\times6.022\times10^{23}}{131}=4.60\times10^{15}"

b) The activity of iodine-131 per gram is known to be 125,000 curie; thus, the activity of 1 mg is "\frac{125,000}{10^6}=0.125" curie

c) 50 days is "\frac{50}{8.02}=6.23" half-life periods

The activity will be "\frac{0.125}{2^{6.23}}=0.00167" curie

d)

"log\frac{a}{a_0}=-\frac{ln2}{t_{1/2}}t=-10^{-6}t \\
t=-\frac{log\frac{a}{a_0}}{10^{-6}} \\
log\frac{0.0001}{0.125}=-3.097 \\
t=\frac{3.097}{10^-6}=3.097\times10^6 \ s" or "3.097\times24\times3600" = 35.8 days